A skier traveling 11.0 m/s reaches the foot of a steady upward 19° incline and

glides 15 m up along this slope before coming to rest. What was the average

coefficient of friction? (use conservation of energy method or work energy

theorem)


1
Expert's answer
2022-05-30T15:17:21-0400
"\Delta E=W"

"0-\frac{mv^2}{2}=-mgl \sin\theta-F_fl"

"F_f=\frac{mv^2}{2l}-mg \sin\theta=\mu mg\cos\theta"

"\mu=\frac{v^2}{2gl\cos\theta}-\tan\theta"

"\mu=\frac{11.0^2}{2*9.8*15\cos19^\circ}-\tan19^\circ=0.091"